OpenStudy (anonymous):

Why does (sqrt2/2)+(sqrt2/2) = sqrt2?

OpenStudy (saifoo.khan):

\[\frac{\sqrt2}{2} + \frac{\sqrt2}{2}\]Take LCD = 2 \[\frac{\sqrt2 + \sqrt2}{2}\]\[\frac{2 \sqrt2 }{2}\]

OpenStudy (mertsj):

What does 1 cat + 1 cat = 2 cats?

OpenStudy (saifoo.khan):

\[\frac{\cancel2 \sqrt2 }{\cancel2} \to \therefore \sqrt2 \]

OpenStudy (saifoo.khan):

1 cat + 1 cat = 1(cat + cat) = 1(2cat) = 2cat

OpenStudy (anonymous):

Thanks. That was helpful.

OpenStudy (saifoo.khan):

Welcome.